OpenStudy (anonymous):

The students at Midtown Middle school sold flowers as a fundraiser in September and October. In October, they charged $1.50 for each flower. The October price was a 20% increase of the September price. Part A What was the price of the flowers in September? Enter your answer in the box. $ Part B The seventh-grade class earned 40% of the selling price of each flower. In September, they sold 900 flowers. In October, they sold 700 flowers. Select a choice from each drop-down menu to make a true statement. The seventh-grade class earned more money in by

OpenStudy (anonymous):

help pls

OpenStudy (anonymous):

do you know percentage already?

OpenStudy (anonymous):

noi need help bro pls

OpenStudy (anonymous):

ok when we say 20%, it means 20 per 100 or 20/100 or 0.2...

OpenStudy (anonymous):

oh ok do u know the answer bro pls

OpenStudy (anonymous):

ok for Part A... if P is the price for September, $1.50 for October, it is the 20% increase of P in September, therefore we solve for P, the equation which I want you to solve is \[P+0.2P=1.50\]

OpenStudy (anonymous):

i dont know how to do that

OpenStudy (anonymous):

thats why im asking for help can u help me bro

OpenStudy (anonymous):

@Orion1213

OpenStudy (anonymous):

all you have to do is perform the arithmetic operations on like terms such as P, what is 1 + 0.2?

OpenStudy (anonymous):

ok let me try

OpenStudy (anonymous):

1.2

OpenStudy (anonymous):

@Orion1213

OpenStudy (anonymous):

so you'll have\[1.2\times P=1.50\]now what is the reverse of multiplication, that is the operation for you to solve P... you get 1.50 when you 1.2 to a number P... so how will you solve P?

OpenStudy (anonymous):

division

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so how will you do it?

OpenStudy (anonymous):

1.2xp=1.5

OpenStudy (anonymous):

i think thats how

OpenStudy (anonymous):

@Orion1213

OpenStudy (anonymous):

solve for p, you said division....

OpenStudy (anonymous):

0.8

OpenStudy (anonymous):

i'll give you a pattern, 2 x 3 = 6, if i replace M for 3, it will be 2 x M = 6, how will i get M = 3? we get 3 when we divide 6 by 2, and so M = 6/2 = 3

OpenStudy (anonymous):

@Orion1213

OpenStudy (anonymous):

i solved for p and got 0.8 @Orion1213

OpenStudy (anonymous):

it seems not... how did you perform it...

OpenStudy (anonymous):

see the pattern i gave...

OpenStudy (anonymous):

1.2 divided by p =1.50

OpenStudy (anonymous):

thats what i did

OpenStudy (anonymous):

@Orion1213

OpenStudy (anonymous):

amd got 0.8

OpenStudy (anonymous):

could it be \[P=\frac{1.5}{1.2}\]

OpenStudy (anonymous):

so P = $1.25 for September, 0.25 is the 20% of it... so if we increase 20% by October, it will be 1.25 + 0.25 = 1.50... hope you understand it...

OpenStudy (anonymous):

ok so whats the answer

OpenStudy (anonymous):

haven't you read between the lines... my last reply?

OpenStudy (anonymous):

we're solving for P...

OpenStudy (anonymous):

im sorry im just really confused

OpenStudy (anonymous):

this is new to me i just moved to us and it hard sorry bro

OpenStudy (anonymous):

all you need is to put an interest in the subject...

OpenStudy (anonymous):

'cause if you don't have any interest to learn MATH, you wouldn't learn MATH at all... even a simple arithmetic operation will be hard for you....

OpenStudy (anonymous):

in my country they teach completly diffrent